### v2

# The kinematic cosmic dipole beyond Ellis and Baldwin

11institutetext: Sorbonne Université, CNRS, Institut d’Astrophysique de Paris, 98bis Boulevard Arago, 75014 Paris, France
The kinematic cosmic dipole beyond Ellis and Baldwin
Albert Bonnefous, albert.bonnefous@iap.fr
(Received $Y$ 20ZZ)
[…]
1 Introduction
In their 1984 article (Ellis and Baldwin 1984), Ellis & Baldwin pointed out that our motion with respect to the Cosmological Rest Frame (CRF) induces an intrinsic anisotropy in the observed sky distribution of any class of objects in the Universe. In other words, the Universe does not appear perfectly homogeneous and isotropic to us. In the specific case of radio sources, whose spectra and flux distributions can be well approximated by power laws, they proposed a method to infer our velocity in the CRF by measuring this anisotropy, more precisely the dipole in the number of sources per unit solid angle. They derived what we will hereafter refer to as the EB formula:
$\bm{\mathcal{D}}_{\rm kin}=(2+x(1+\alpha))\bm{\beta}\,$, (1)
where $x$ and $\alpha$ denote the power-law indices of the flux distribution and source spectra, respectively, and $\bm{\beta}$ is our velocity relative to the CRF, normalized by the speed of light. Ellis & Baldwin argued that estimating $\bm{\beta}$ in this way should yield the same result as the velocity measured from the dipole of the Cosmic Microwave Background (CMB), provided the sources are sufficiently distant that intrinsic clustering anisotropies can be neglected, and lie entirely beyond our local bulk flow, in a Friedmann-Lemaître-Robertson-Walker (FLRW) Universe.
[…]
These findings have renewed interest in the EB formalism and motivated efforts to extend its application to upcoming surveys such as Euclid (Euclid Collaboration et al. 2025), LSST (Rubin Observatory) (Ivezić et al. 2019), and SPHEREx (Crill et al. 2020). In particular, the expression for the kinematic dipole in number counts has been generalized to a tomographic dipole that incorporates redshift evolution in the source description (Maartens et al. 2018), with alternative expression involving the time evolution of sources (Dalang and Bonvin 2022). It has also been shown that the original EB formula remains correct in the presence of source evolution, and that in the case of monochromatic surveys it is necessary to evaluate the parameters $x$ and $\alpha$ at the flux limit (von Hausegger 2024). A more detailed review on this test and its use through the year is available in (Secrest et al. 2025). However, the case of sources that do not exhibit power law spectra in photometric survey, has not been treated yet. For example, the spectra of galaxies in the visible and near infrared can feature emission line, bumps, or even more complicated features such as Lyman-$\alpha$ forest, see Fig. 1. Solving this issue is crucial to apply this test to upcoming large scale surveys such as LSST (Ivezić et al. 2019) or Euclid (Euclid Collaboration et al. 2025).
In this letter, we first briefly review the formalism underlying the EB formula in Section 2. We then extend it in Section 3 to the general case of a source population without assuming any specific spectral profile or luminosity distribution, considering two distinct survey types: monochromatic (measuring the flux density, as is typical of radio surveys) and photometric (the flux is integrated with a pass-band filter). Finally, in Section 4, we apply the resulting effective spectral index in photometric surveys to quasars and compare it to the value obtained using the approach of (Secrest et al. 2021).
Figure 1: Sketch of an integral number count (left) with a power law fit in dashed red at an arbitrary flux or luminosity limit, and an arbitrary spectra (right) of a source at the flux or luminosity limit, where the red part correspond to the part of the spectra integrated in a given band $X$. In the most general case survey, it is always possible to fit a power law at the flux or luminosity limit. However here the spectra is not a power law, so in the case of a photometric survey the spectral index $\alpha$ cannot be defined as $S_{\nu}\propto\nu^{-\alpha}$ over the band.
2 The Ellis & Baldwin formula
Consider a moving observer with respect to a rest frame, such as presented in Fig. 2. Any quantity $a$ measured in the frame of the observer will be written $a_{\rm o}$, whereas the same quantity in the rest frame will be $a_{\rm r}$. Two different effects impact the observations, the relativistic aberrations, and the Doppler boosting.
Figure 2: (a) The frame moving with the observer, (b) the rest frame, in which the observer has a velocity $\mathbf{v}$. For the sake of simplicity, we assume that the $z$ axis is parallel to the movement of this observer.
2.1 Relativistic aberration
The relativistic aberration of light is the deformation of the field of view of an observer moving with respect to what it observes; the apparent position of a body is shifted toward the direction of its movement. Suppose that we are looking at an object at a position $\bm{\theta}_{\rm o}=(\theta_{\rm o},\phi_{\rm o})$ in the sky. In the rest frame, the angle $\phi_{\rm r}$ does not change, and $\theta_{\rm r}$ transforms such that:
$\sin{\theta_{\rm r}}=\frac{\sin{\theta_{\rm o}}}{\gamma(1-\beta\cos{\theta_{\rm o}})}\,$, (2)
where $\gamma=1/(1-\beta^{2})^{1/2}$. With ${\rm d}\Omega=\sin{\theta}{\rm d}\theta{\rm d}\phi$, we obtain the distortion of the solid angles relation:
${\rm d}\Omega_{\rm o}=\delta(\bm{\theta}_{\rm o})^{-2}{\rm d}\Omega_{\rm r}\,$, (3)
and introducing the notation:
$\delta(\bm{\theta})=\frac{1}{\gamma(1-\bm{\beta}\cdot\bm{\theta})}\approx 1+\bm{\beta}\cdot\bm{\theta}+\mathcal{O}(\beta^{2})\,.$
Also, for the sake of simplicity, the $\mathcal{O}(\beta^{2})$ will be omitted in every approximation that follows.
2.2 Relativistic Doppler effect
The Doppler boosting effects describes that the light coming from a distant object will be brighter, and bluer, in the direction of the movement of the observer. Between the two frames, the frequency of a given photon shifts as:
$\nu_{\rm o}=\delta(\bm{\theta}_{\rm o})\nu_{\rm r}\,.$ (4)
The flux density $S=\frac{{\rm d}E}{{\rm d}\nu{\rm d}t}$ also transforms between frames as:
$S_{\rm o}(\nu_{\rm o})=\delta(\bm{\theta}_{\rm o})S_{\rm r}(\nu_{\rm r})\,.$ (5)
2.3 Ellis & Baldwin formula with power laws
Now, let us assume that the rest frame is the CRF. If the observer had no movement with respect to this frame, they would observe a completely uniform background of NN objects, this uniformity provided that they are sufficiently far away that local inhomogeneities can be negligible (Gibelyou and Huterer 2012). These objects can be galaxies, quasars, etc., but we will consider that there is only one type of object in our data. Ellis & Baldwin (Ellis and Baldwin 1984) make the following assumptions. First, every object’s spectral flux density $S$ has a simple power law dependence on the frequency $\nu$:
$S\propto\nu^{-\alpha}\,.$ (6)
Secondly, the objects have a power law distribution in the number of objects NN above a certain flux $S_{\rm lim}$:
$N(>S_{\rm lim})\propto S_{\rm lim}^{-x}\,.$ (7)
These two indices $x$ and $\alpha$ completely depend on the type of object we observe, their redshift distribution and the chosen flux cut. In the observer’s frame, the flux received from one source is boosted by the Doppler effect, and using the power law (6), the flux received from this object at a fixed frequency $\nu_{\rm o}$ is:
$S_{\rm o}(\nu_{\rm o})=\delta(\bm{\theta}_{\rm o})^{1+\alpha}S_{\rm r}(\nu_{\rm o})\,.$ (8)
The number of objects, in a given direction $\bm{\theta}_{\rm o}$, then changes as:
$\begin{split}{\rm d}N_{\rm o}(>S_{\rm lim},\bm{\theta}_{\rm o})&={\rm d}N_{\rm r}(>\delta(\bm{\theta}_{\rm o})^{-(1+\alpha)}S_{\rm lim},\bm{\theta}_{\rm o})\\ &=\delta(\bm{\theta}_{\rm o})^{x(1+\alpha)}{\rm d}N_{\rm r}(>S_{\rm lim},\bm{\theta}_{\rm o})\,.\end{split}$ (9)
Note that here, $S_{\rm lim}$ is fixed and it is the objects’ flux density that changes, which is as if the flux density limit depended on the frame. Finally, taking into account the aberration of the solid angles (3), the number count of objects by solid angle $N_{\Omega}=\frac{{\rm d}N}{{\rm d}\Omega}$ in the observer frame becomes:
$\begin{split}N_{\Omega,{\rm o}}(>S_{\rm lim},\bm{\theta}_{\rm o})&=\frac{{\rm d}N_{\rm o}}{{\rm d}\Omega_{\rm o}}(>S_{\rm lim},\bm{\theta}_{\rm o})\\ &=\delta(\bm{\theta}_{\rm o})^{2+x(1+\alpha)}N_{\Omega,{\rm r}}(>S_{\rm lim})\,.\end{split}$ (10)
If we integrate this formula on the whole sky, the overall number of observed sources in the sky does not change with the frame at order $\beta$, and we can write
$N_{\Omega,{\rm r}}(>S_{\rm lim})\approx\bar{N}_{\Omega}(>S_{\rm lim})\approx N(>S_{\rm lim})/4\pi\,.$
Note that if we want to take into account any other non-kinematic cosmological perturbation in the number count, we can do so by including them in this $\bar{N}_{\Omega}$ (Nadolny et al. 2021). Now, if we take the CMB as a reference frame, we have $v=369.82\pm 0.11\,{\rm km.s^{-1}}$ (Collaboration et al. 2020). Therefore $\beta\ll 1$, and the number count (10) can be expressed purely with observed quantities:
$N_{\Omega,{\rm o}}(>S_{\rm lim},\bm{\theta}_{\rm o})\approx\bar{N}_{\Omega}(1+\bm{\mathcal{D}}_{\rm kin}\cdot\bm{\theta}_{\rm o})$ (11)
$\bm{\mathcal{D}}_{\rm kin}=(2+x(1+\alpha))\bm{\beta}$ (12)
This is the form derived by Ellis & Baldwin (Ellis and Baldwin 1984). We’ve seen that this expressions assumes that the spectrum and luminosity function of these objects are power laws, and most of all that there exists a cosmological rest frame, in which there is a uniform background of these objects. However, since aberration and Doppler boosting are purely local relativistic effects, this formula is completely independent of the cosmological model provided that the cosmological principle holds. This formula also requires no knowledge of the redshift of the objects, only their location in the sky.
3 Kinematic dipole without any assumption on the spectrum and number count
Here, we want to derive a dipole expression, without making the power law hypothesis for the cumulative number count and spectrum of our sources. The assumptions we will make is that $\beta\ll 1$ and that we observe the same spectra for all of the sources, or at least the same spectral profile. In other words $S_{i}\propto s(\nu)$ for every object $i$ in our survey, and $s(\nu)$ any function.
3.1 Monochromatic survey
First, we suppose that we are looking at a survey at a single wavelength $\nu_{\rm o}$. At this wavelength, the spectral flux density $S$ transforms as:
$\begin{split}S_{\rm o}(\nu_{\rm o})&=\delta(\bm{\theta}_{\rm o})S_{\rm r}(\nu_{\rm o}\delta(\bm{\theta}_{\rm o})^{-1})\\ &\approx(1+\bm{\beta}\cdot\bm{\theta}_{\rm o})\left(S_{\rm r}(\nu_{\rm o})-\nu_{\rm o}\frac{\partial S_{\rm r}}{\partial\nu_{\rm o}}\bm{\beta}\cdot\bm{\theta}_{\rm o}\right)\,.\end{split}$ (13)
We obtain $S_{\rm o}(\bm{\theta}_{\rm o},\nu_{\rm o})=S_{\rm r}(\nu_{\rm o})+\delta S(\bm{\theta}_{\rm o},\nu_{\rm o})$, where $\delta S(\bm{\theta}_{\rm o},\nu_{\rm o})$ is defined as:
$\begin{split}\delta S(\bm{\theta}_{\rm o},\nu_{\rm r})&\approx\left(S_{\rm r}(\nu_{\rm o})-\nu_{\rm o}\frac{\partial S_{\rm r}}{\partial\nu_{\rm o}}\right)\bm{\beta}\cdot\bm{\theta}_{\rm o}\\ &\approx S(\nu)\left(1-\frac{\partial\log S}{\partial\log\nu}\right)\bm{\beta}\cdot\bm{\theta}_{\rm o}\,.\end{split}$ (14)
Note that here, every quantity in the rest frame can be expressed in the observer frame since the difference is in $\mathcal{O}(\beta)$, and becomes $\mathcal{O}(\beta^{2})$ after multiplication by $\bm{\beta}\cdot\bm{\theta}_{\rm o}$. Therefore, for the sake of simplicity in the following we will drop the indices indicating the frame for every term proportional to $\bm{\beta}$. The number count becomes:
$\begin{split}&N_{\Omega,{\rm o}}(>S_{\rm lim},\bm{\theta}_{\rm o})=\delta(\bm{\theta}_{\rm o})^{2}N_{\Omega,{\rm r}}(>S_{\rm lim}-\delta S_{\rm lim},\bm{\theta}_{\rm r})\\ &\approx(1+2\bm{\beta}\cdot\bm{\theta}_{\rm o})\left(\bar{N}_{\Omega}(>S_{\rm lim})-\frac{\partial\bar{N}_{\Omega}}{\partial S_{\rm lim}}\delta S_{\rm lim}\right)\\ &\approx\bar{N}_{\Omega}\left(1+\left(2-\frac{\partial\log N}{\partial\log S_{\rm lim}}\left(1-\frac{\partial\log S_{\rm lim}}{\partial\log\nu}\right)\right)\bm{\beta}\cdot\bm{\theta}_{\rm o}\right)\,.\end{split}$ (15)
As before, in the rest frame, $N_{\Omega,{\rm r}}(>S_{\rm lim,{\rm r}},\bm{\theta}_{\rm r})\approx N(>S_{\rm lim})/4\pi$, and the kinematic dipole in the observer frame is:
$\bm{\mathcal{D}}_{\rm kin}=\left(2-\frac{\partial\log N}{\partial\log S_{\rm lim}}\left(1-\frac{\partial\log S_{\rm lim}}{\partial\log\nu}\right)\right)\bm{\beta}\,.$ (16)
This expression is strictly equivalent to the EB formula (12), where we took effective coefficient:
$x_{\rm eff}=-\frac{\partial\log N}{\partial\log S_{\rm lim}}$ (17)
at the limiting flux density $S_{\rm lim}$. This is expected since only the objects with a flux density close to the limit appear or disappear from the number count with Doppler boosting.
$\alpha_{\rm eff}=-\frac{\partial\log S_{\rm lim}}{\partial\log\nu}$ (18)
at the frequency $\nu_{\rm o}$. These results are similar as the one obtained in (von Hausegger 2024), noting that here since we considered that all of our sources have the same spectral profile, this $\alpha_{\rm eff}$ doesn’t have to be expressed at $S_{\rm lim}$ specifically, but in the general case this spectral index has to be taken for the sources at the flux limit, which are the one susceptible to appear or disappear with Doppler boosting. In particular, it is notable that when the observed wavelength is close to a strong emission line, and every source have the same redshift, this $\alpha_{\rm eff}$ could become huge, in the negative or positive depending on the position of this emission line with $\nu_{\rm o}$. This could lead to a sudden jump in the number count dipole. However, an important property of the spectral index is lost in this general case, it is no longer redshift independent. If we take into account that the spectra of every sources get shifted according to their individual redshift $S_{i}(\nu)\rightarrow S_{i}^{\rm int}((1+z_{i})\nu)$, where $S_{i}^{\rm int}$ is the intrinsic spectra of the sources, the $\alpha_{\rm eff}$ also gets modified. However, this doesn’t affect the power law, as the redshift doesn’t impact the spectrum profile, and it stays a power law.
3.2 Photometric survey
In practice, cosmological surveys often don’t scan the sky at a monochromatic frequency $\nu_{\rm o}$, but give the luminosity (or equivalently, the magnitude) of every objects in a given band $X$. This luminosity can be expressed using the transmission function of the filter $T_{X}(\nu)$, here in terms of received energy:
$L=\int{\rm d}\nu\,T_{X}(\nu)S(\nu)\,.$ (19)
Note that here we neglected the transmission function of the atmosphere itself for ground based telescope, but the effect of atmosphere can be included in this filter transmission function. Suppose that we are looking at a particular object in the direction $\bm{\theta}_{\rm o}$ whose spectral flux density is $S_{\rm o}(\nu_{\rm o})$, its luminosity transforms in the observer frame as:
$\begin{split}&L_{\rm o}(\bm{\theta}_{\rm o})=\int{\rm d}\nu_{\rm o}\,T_{X}(\nu_{\rm o})S_{\rm o}(\nu_{\rm o})\\ &=\delta(\bm{\theta}_{\rm o})^{2}\int{\rm d}\nu_{\rm r}\,T_{X}(\delta(\bm{\theta}_{\rm o})\nu_{\rm r})S_{\rm r}(\nu_{\rm r})\\ &\approx(1+2\bm{\beta}\cdot\bm{\theta}_{\rm o})\int{\rm d}\nu_{\rm r}\,\left(T_{X}(\nu_{\rm r})+\bm{\beta}\cdot\bm{\theta}_{\rm o}\nu_{\rm r}\frac{\partial T_{X}}{\partial\nu_{\rm r}}\right)S_{\rm r}(\nu_{\rm r})\,.\end{split}$ (20)
We obtain $L_{\rm o}(\bm{\theta}_{\rm o})=L_{\rm r}+\delta L(\bm{\theta}_{\rm o})$, where $\delta L(\bm{\theta}_{\rm o})$ is defined as:
$\begin{split}\delta L(\bm{\theta}_{\rm o})&\approx\left(2L+\int{\rm d}\nu\,\nu\frac{\partial T_{X}}{\partial\nu}S(\nu)\right)\bm{\beta}\cdot\bm{\theta}_{\rm o}\\ &\approx\left(L-\int{\rm d}\nu\,T_{X}(\nu)\nu\frac{\partial S}{\partial\nu}\right)\bm{\beta}\cdot\bm{\theta}_{\rm o}\,.\end{split}$ (21)
Note that we used $T_{X}(\nu)=T_{X}^{\prime}(\nu)=0$ outside of a particular frequency interval. For the same reason as $\delta S$ in the monochromatic survey, every quantity here can be expressed either in the rest frame or in the observer frame. The number count becomes:
[…]
where $S_{\rm lim}(\nu)$ here is such as that its integrated luminosity is $L_{\rm lim}=\int{\rm d}\nu\,T_{X}(\nu)S_{\rm lim}(\nu)$. As before, we can define effective coefficient $x_{\rm eff}$ and $\alpha_{\rm eff}$ where:
$x_{\rm eff}=-\frac{\partial\log N}{\partial\log L_{\rm lim}}=+2.5\frac{\Phi(m_{\rm lim})}{N}\,$, (23)
with $\Phi(m)=\frac{{\rm d}N}{{\rm d}m}$ the luminosity function of the survey, $m_{\rm lim}$ the magnitude corresponding to $L_{\rm lim}$, and:
$\alpha_{\rm eff}=-\frac{\int{\rm d}\nu\,T_{X}(\nu)\nu\frac{\partial S_{\rm lim}}{\partial\nu}}{\int{\rm d}\nu\,T_{X}(\nu)S_{\rm lim}(\nu)}=1+\frac{\int{\rm d}\nu\,\frac{\partial T_{X}(\nu)}{\partial\nu}\nu S_{\rm lim}(\nu)}{\int{\rm d}\nu\,T_{X}(\nu)S_{\rm lim}(\nu)}\,.$ (24)
The precedent remark on the necessity to take this spectral index at the flux limit in the general case is still valid here, even though within our simplifying assumption that all sources have the same spectral profile it can be simplified. Note that with $T_{X}(\nu)=\nu_{\rm o}\delta(\nu-\nu_{\rm o})$ for a given wavelength $\nu_{\rm o}$, we obtain the same coefficients as for the monochromatic survey, which is expected. If we also suppose that $S\propto s(\nu)=\nu^{-\alpha}$, we have $\alpha_{\rm eff}=\alpha$, whatever be the filter $T_{X}(\nu)$. We have shown that the EB formula (12) is still valid for any spectrum profile and luminosity function in photometric survey, provided that the $x$ and $\alpha$ coefficient are defined with more precision. However, as before we lost the property of the redshift independence of this spectral index.
4 Effective spectral index for quasars
Now that we have an expression for the general effective coefficient for any spectral profile, we want to apply this new expression of $\alpha_{\rm eff}$ to real data, as a proof of concept, in order to probe the limitations of this formula. In particular, we want to see if the method to obtain $\alpha$ coefficient used in (Secrest et al. 2021), which we will denote $\alpha_{\rm mag}$ in the following, yields the same results as this one. Using the W1 and W2 bands of the CatWISE survey, this method compares the difference of magnitude W1-W2, with a table of the same quantity calculated for synthetic pure-power law spectra with different $\alpha$ coefficient.
To obtain $\alpha_{\rm eff}$ coefficients, we need quasar spectra in this W1 band, which expands from $2.7$ to $3.9{\rm\penalty 10000\ \mu m}$ (Eisenhardt et al. 2020). We will use the AKARI QSONG catalog (Myungshin et al. 2017), which is a compilation of $753$ quasars’ spectra taken between $2.5$ and $5{\rm\penalty 10000\ \mu m}$. However, since we want a good spectral quality, we will only keep the spectra with a mean spectral flux density above $5{\rm\penalty 10000\ mJy}$ and a positive value for $\alpha_{\rm fit}$, which brings 41 individual quasars, and we can then extract their corresponding W1 and W2 band magnitude from the CatWISE2020 catalog (Marocco et al. 2021). We also use the transmission filter $T_{\rm W1}(\nu)$, given by CatWISE  [^1] The transmission in response per photon is given here https://www.astro.ucla.edu/~wright/WISE/passbands.html.
Now, we can apply equation (24) to these spectra to obtain the effective spectral index $\alpha_{\rm eff}$. We will use the Python package uncertainy to estimate the error bars  [^2] See https://pythonhosted.org/uncertainties/. It has to be noted that a naive integration of (24) can lead to huge error bars in the result: the gradient of the spectra is often smaller than its uncertainty here; however the gradient is not a set of independent variables, it is constrained by the value of the sum of it, and therefore the uncertainty in this gradient can be easily over-estimated. Error propagation must be carefully handled, and it may be better suited to use the last term of the expression (24) which doesn’t involve this gradient. Then, we can infer the corresponding spectral index $\alpha_{\rm mag}$, with the CatWISE magnitude. Lastly, we will also fit the spectra within the range of the W1 band with a power law to obtain a last spectral index $\alpha_{\rm fit}$.
Figure 3: Comparison of the spectral index $\alpha$ obtained with different methods, for 41 quasar’s spectra. The black dashed line corresponds to the identity. The two red dots with a high deviation between their calculated $\alpha_{\rm eff}$ and the two other method corresponds to the two quasars with the lowest mean spectral flux density.
We compare the different spectral indices obtained with these three methods in Fig. 3. We can see that the uncertainty in calculating $\alpha_{\rm eff}$ is quite large, which is explained by the large uncertainty of the spectra themselves. It is notable that the two point with the biggest deviation between this $\alpha_{\rm eff}$ and the other two methods correspond to the two spectra with the lowest spectral flux density. This emphasize the necessity to have a good spectral quality to calculate $\alpha_{\rm eff}$ precisely. For all of these quasars, we obtain:
[…]
We lack the statistical significance to make any definitive claim, however nothing leads us to think that these three methods yields significantly different results for quasars in this $\rm W1$} band. At most, it seems that the $\alpha_{\rm mag}$ are slightly smaller than the other spectral indexes. If we consider that this method underestimate by $0.16$ the $\alpha$, and with a mean measured value for the quasars in (Secrest et al. 2021) for $\alpha$ being $1.26$, the quantity $2+x(1+\alpha)$, with $x=1.89$, would be at most underestimated by about $4\%$. This discrepancy might be explained by a slight difference in slope between the $\rm W1$} and $\rm W2$} band, but further investigation would be necessary to obtain conclusive evidences. Lastly, an example of a quasar spectrum taken from AKARI with its corresponding $\alpha_{\rm eff}$ and $\alpha_{\rm mag}$ is shown in Fig. 4.
Figure 4: Example of a spectrum for a quasar (PG2112+059, located at $(272.101907,64.128918)$, in galactic coordinates) given by AKARI, with the spectral indexes $\alpha_{\rm mag}$ obtained using the $\rm W1-W2$} method, and $\alpha_{\rm eff}$ using a direct integration of the spectra using (24). These spectral indexes are plotted as pure power laws $S_{\nu}\propto\nu^{-\alpha}$.
5 Conclusion
In this letter, we have demonstrated that the Ellis & Baldwin formula can be generalized to any class of light sources, regardless of their spectral shape or luminosity distribution, provided the coefficients $x$ and $\alpha$ are appropriately redefined. This result is essential for addressing the cosmic dipole anomaly, as it shows that sources with more complex spectral profile than power laws such as galaxies in the near-infrared or visible light can in principle be used to measure our velocity relative to the CRF with photometric surveys. Upcoming large scale photometric surveys such as LSST (Ivezić et al. 2019) or Euclid (Euclid Collaboration et al. 2025) will provide us with the opportunity to apply the Ellis and Baldwin test to such data.
We have also shown that, in the specific case of quasars, using spectra from AKARI, the resulting effective coefficient $\alpha$ within the CatWISE $\rm W1$} band is generally consistent with the value obtained by fitting the spectrum with a pure power law. Moreover, we find that the procedure used in (Secrest et al. 2021), which relies only on the $\rm W1$} and $\rm W2$} magnitudes, may only slightly underestimate the true spectral index by up to $\sim 0.16$. Still, this difference is not sufficient to significantly alter the conclusions of this work, that has shown that the quasar dipole is a factor of $\sim 2$ larger than the kinematic dipole expectation. Nevertheless, the limited number of available quasar spectra in this band prevents us from drawing definitive conclusions on this point. In the future, this issue may be clarified using spectroscopic libraries from missions such as SPHEREx (Crill et al. 2020).
In the general case, however, determining this effective spectral index $\alpha_{\rm eff}$ rigorously for a given dataset remains challenging. First, the spectral index is no longer independent of redshift. Beyond this, another important issue is that most wide-area sky surveys are photometric, and the calculation of $\alpha_{\rm eff}$ requires a good quality of spectra within the range of the photometric band. Spectroscopic surveys have generally much less objects and this index may only be available for only a small subset of sources through cross-matching. Good quality spectra is generally only available for more luminous objects, therefore measuring $\alpha_{\rm eff}$ for sources at the flux limit specifically would be complicated. This may introduce biases which will have to be carefully handled. Future applications of this test will need to address these issues directly.
Acknowledgements.
The author thanks Roya Mohayaee, Nathan Secrest, Reza Ansari, Johann Cohen-Tanugi, Reiko Nakajima and Sebastian Von Hausegger for their important insights and help.